Balance equation for aluminum acetate and ammonium phosphate. If there is 12.579 g of aluminum acetate and 10.326 g of ammonium phosphate what is the theoretical yield? What is the limiting reagent?
1
Expert's answer
2012-10-23T09:38:52-0400
(CH3COO)3Al + (NH4)3PO4 = 3AlPO4 + 3CH3COONH4

n of
(CH3COO)3Al =12.579/204.114=0.06163 mol

n of (NH4)3PO4 = 10.326/149.087=
0.06926 mol ( (NH4)3PO4 in excess,so (CH3COO)3Al is limiting reagent )